# Federated Hierarchical Reinforcement Learning for Adaptive Traffic Signal Control

### v1

@@ 3.2.1 Client Side
In FRL, each client $n\in\mathcal{N}=\{1,\dots,N\}$ corresponds to a traffic light with local model parameters $w_{n}^{\tau}$ in the local round $\tau\in\{1,\dots,\mathcal{T}\}$. At the start of each communication round, a central FL server distributes the current global model to a subset of these clients. Each client then trains its local model by optimizing a local reinforcement learning objective $R_{n}(w_{n}^{\tau})$, using data gathered from its own environment. Once training is complete, each client uploads its updated parameters to the server. The server aggregates these updates and redistributes the resulting global model for the next round. This process facilitates collaborative learning across multiple agents, reduces communication overhead, and preserves data privacy. In its simplest form, each agent’s parameters may be updated as follows:
@@
Each agent interacts with its local environment for one or more episodes or time steps. During this phase, it collects trajectories of states, actions, and rewards:
@@
Here, $s_{n}^{\tau}$ is the traffic state of agent $n$ at time step $\tau$, $a_{n}^{\tau}$ is the selected action, and $R_{n}^{\tau}$ is the corresponding reward.
@@ 4.2 FedFomoLight
We adapt the general FedFomo framework to MARL, resulting in the proposed FedFomoLight. Each client $n$ maintains a personalized objective aimed at maximizing its local RL reward $\mathcal{R}_{n}(w^{t}_{n})$ given the current local model parameters $w^{t}_{n}$ in round $t$. The global optimization objective is then defined as finding the optimal set of local model parameters:
$\bigl{\{}(w^{t}_{1})^{*},(w^{t}_{2})^{*}\dots,(w^{t}_{|N|})^{*}\bigr{\}}\;=\;\arg\!\max_{\{w^{t}_{n}\}_{n=1}^{|N|}}\;\sum_{n=1}^{|N|}\mathcal{R}_{n}(w^{t}_{n}),$ (7)
Thus, each agent strives to maximize its personalized RL objective $\mathcal{R}_{n}$, while still benefiting from collaborative updates provided by the federated framework. Formally, this MARL-adapted formulation parallels the supervised definition in the original FedFomo framework, but incorporates RL-specific objectives $\mathcal{R}_{n}$.
@@
Unlike FedAvg, which assigns equal weight to all models (or weights them by local dataset size), for each agent $n$ FedFomoLight optimizes personalized importance $\boldsymbol{\rho_{n}}=[\rho_{n1},\dots,\rho_{nN}]$, where $\sum_{i=1}^{N}\rho_{ni}=1$ to improve the local RL objective $\mathcal{R}_{n}$. We do this by updating $\rho_{ni}$ via gradient descent:
[…]
@@
[…]
To approximate $\frac{\partial}{\partial\rho_{ni}}\,\mathcal{R}_{n}(w_{n}^{t-1})$, we employ a first-order Taylor expansion. Define:
@@
Evaluating $\mathcal{R}_{n}(\varphi_{ni}(\rho_{ni}))$ near $\rho_{ni}=0$ and then setting $\rho_{ni}=1$ provides:
$\frac{\partial}{\partial\rho_{ni}}\,\mathcal{R}_{n}\bigl{(}w_{n}^{t-1}\bigr{)}\;\approx\;\mathcal{R}_{n}\bigl{(}w_{i}^{t}\bigr{)}\;-\;\mathcal{R}_{n}\bigl{(}w_{n}^{t-1}\bigr{)}.$ (12)
@@
$\rho_{ni}\;=\;-\,\alpha\,\frac{\mathcal{R}_{n}\bigl{(}w_{i}^{t}\bigr{)}\;-\;\mathcal{R}_{n}\bigl{(}w_{n}^{t-1}\bigr{)}}{\bigl{\|}w_{i}^{t}-w_{n}^{t-1}\bigr{\|}}.$ (13)
@@
Algorithm 1 FedFomoLight
1:Initialize: For each agent $n\in\{1,2,\cdots,N\}$, initialize local parameters $w_{n}^{0}$
2:Set: Number of global rounds $T$, local training steps per round $\mathcal{T}$
3:for each global round $t=1,2,\cdots,T$ do
@@
6:         Run A2C locally for $\mathcal{T}$ timesteps using $w_{n}^{t-1}$
7:         Obtain updated local parameters $w_{n}^{t}$
8:         Upload $w_{n}^{t}$ to the central FL server
@@
13:         Let $\boldsymbol{\rho}_{n}=[\rho_{n1},\rho_{n2},\cdots,\rho_{nN}]$ be the importance vector for agent $n$
14:         for each agent $i\in\{1,2,\cdots,N\}$ do
15:              Compute raw importance $\rho_{ni}$ via Equation (13)
@@
17:         Normalize the importance vector to ensure $\rho_{ni}\geq 0$ and $\sum_{i}\rho_{ni}=1$
@@
21:         Update model parameters using Equ. (15)
22:         Return updated $w_{n}^{t}$ to agent $n$
@@ 4.3 FedClusterLight
1:Initialize: For each agent $n\in\{1,2,\cdots,N\}$, initialize local Actor and Critic parameters $w_{n}^{0}$
2:for each global round $t=1,2,\cdots,T$ do
@@
5:         Run A2C locally for $\mathcal{T}$ timesteps using $w_{n}^{t-1}$
6:         Obtain updated local parameters $w_{n}^{t}$
7:         Upload $w_{n}^{t}$ to the central FL server
@@
10:     The server groups agents into $K$ clusters $\{C_{(1)},C_{(2)},\cdots,C_{(K)}\}$ based on similarity of parameters or other relevant attributes
@@
12:     for each cluster $C_{(k)}$ in $\{C_{(1)},C_{(2)},\cdots,C_{(K)}\}$ do
@@
17:         Determine the cluster $k$ such that $n\in C_{(k)}$
18:         Broadcast the weights to the agents according to Equation (18)
@@ 5.1 Experiment Set-up
We utilize Ray’s RLlib to train the proposed HFRL methods and evaluate the final policies via simulation. The road networks are shown in Fig. 3: (1) Grid 3×3 (see Fig. 3) (2) Grid 5×5 (see Fig. 3); (3) Real-world (see Fig. 3). For the 3×3 network, the primary east-west and north-south routes are designed as two-way, four-lane roads, while all other roads are two-way, two-lane roads; for the 5×5 network, the most central north-south and east-west routes are designed as two-way, six-lane roads, adjacent roads are two-way, four-lane roads, and the outermost roads are two-way, two-lane roads. The real-world network is the road networks near Columbia University generated from the OpenStreetMap [^1] https://www.openstreetmap.org, © OpenStreetMap contributors.. For the Grid 3×3 and Grid 5×5 configurations, we designated the rightmost lane at each intersection as a mixed lane for straight-through and right-turn movements, allowing some vehicles to proceed through the intersection or turn right. The straight-through and right-turn ratios for these lanes are set at 90% and 10%, respectively. Regarding the traffic demand, the light demand is $200$ vehicles per lane per hour, while the heavy demand is $400$ vehicles per lane per hour. To build a more realistic simulation on the real-world network, we utilize two data sources to calibrate the simulation: TomTom data from the TomTom MOVE web portal (available at https://move.tomtom.com) and manually recorded traffic counts at intersections. We then used data from the TomTom dataset to define input flows at each edge intersection, as illustrated in Fig. 4. In addition, we collected real-world traffic data from several intersections over a one-hour period, and these data will also be incorporated into our traffic flow generation file.
@@ 5.1.1 Baseline
    Centralized RL. This approach trains a single agent with access to global observations from all intersections. Because the agent sees the entire network state, it can in principle find a coordinated solution, but the communication and computational demands can be high for large networks.
@@ 5.1.2 Performance Metrics
    Travel Time. This refers to the cumulative time each vehicle travels from its origin to its destination within the simulation. Lower travel time usually indicates more efficient traffic flow.
@@
    Waiting Time. This is the total time a vehicle is stopped at intersections throughout its journey. Reducing waiting time reflects improved signal timing and less congestion.
@@
    Communication Cost. This is an estimated communication cost (Hudson et al., 2022) including: (1) transmitting critic and actor parameters from edge devices to the central server; (2) distributing these parameters from the central server back to the edge devices; (3) sending actions from edge devices to traffic lights; (4) relaying observations from traffic lights to edge devices; and (5) transmitting vehicle data from vehicles to the traffic lights.
@@ 5.2.2 Performance Comparison
For the Grid 3×3 network, we present comparisons of travel time and waiting time among our proposed FedClusterLight, FedFomoLight, and baseline methods in Fig. 7 and Fig. 8, respectively. Fig. 7(a) shows that for the Grid 3×3 under relatively low traffic demand, the centralized method outperforms both the decentralized and FedAvg methods. FedClusterLight achieves results close to those of FedAvg, while the proposed FedFomoLight method ranks second only to the centralized approach. According to the results presented in Fig. 7, the decentralized method fails to converge during training, resulting in extremely poor test performance. In contrast, the centralized method achieves the best performance —surpassing FedAvg — while FedFomoLight and FedClusterLight exhibit similar results that are slightly better than FedAvg. As for the waiting time in Fig. 8, the general trend is similar to that of the travel time.
(a) Average travel time on Grid 3×3 under light demand
(b) Average travel time on Grid 3×3 under heavy demand
Figure 7: Average Travel Time on Grid 3×3
(a) Average waiting time on Grid 3×3 under light demand
(b) Average waiting time on Grid 3×3 under heavy demand
Figure 8: Average Waiting Time on Grid 3×3
For the Grid 5×5 network, a similar trend is observed. In light traffic conditions according to Fig. 9(a), the centralized method remains the best, while FedFomoLight and FedClusterLight again outperform FedAvg, with FedFomoLight showing a slight advantage. In heavy traffic, the decentralized performance is significantly worse due to poor convergence in the training process, resulting in the highest travel time. FedFomoLight maintains a superior performance compared to FedAvg and FedClusterLight.
(a) Average travel time on Grid 5×5 under light demand
(b) Average travel time on Grid 5×5 under heavy demand
Figure 9: Average Travel Time on Grid 5×5
(a) Average waiting time on Grid 5×5 under light demand
(b) Average waiting time on Grid 5×5 under heavy demand
Figure 10: Average Waiting Time on Grid 5×5
We compare communication costs across different traffic networks under various traffic demand scenarios in Tab. 4. The centralized method typically requires uploading all local observations to a central server for computation, which leads to significantly higher communication costs. In contrast, FRL and the decentralized method generally maintain similarly low communication costs. FRL methods, due to the allowance for model information exchange, ensure more stable communication costs during training. Furthermore, the decentralized method, limited by partial observations, tends to be less stable in complex traffic scenarios, resulting in higher communication costs.

### v2

@@ 3.2.1 Client Side
In FRL, each client $n\in\mathcal{N}=\{1,\dots,N\}$ corresponds to a traffic light with local model parameters $w_{n}^{\tau}$ in the local round $\tau\in\{1,\dots,\mathcal{T}\}$, where $\mathcal{T}$ denotes the total number of local interaction time steps within one communication round. At the start of each communication round, a central FL server distributes the current global model to a subset of these clients. Each client then trains its local model by optimizing a local reinforcement learning objective $\mathcal{R}_{n}(w_{n}^{\tau})$, using data gathered from its own environment. Once training is complete, each client uploads its updated parameters to the server. The server aggregates these updates and redistributes the resulting global model for the next round. This process facilitates collaborative learning across multiple agents, reduces communication overhead, and preserves data privacy. In its simplest form, each agent’s parameters may be updated as follows:
@@
Each agent interacts with its local environment for $\mathcal{T}$ rounds. During this phase, it collects trajectories of states, actions, and rewards:
@@
Here, $s_{n}^{\tau}$ is the traffic state of agent $n$ at round $\tau$, $a_{n}^{\tau}$ is the selected action, and $R_{n}^{\tau}$ is the corresponding reward.
@@ 4.2 FedFomoLight
We adapt the general FedFomo framework to MARL, resulting in the proposed FedFomoLight. Each client $n$ maintains a personalized objective aimed at maximizing its objective function $\mathcal{L}_{n}(w^{t}_{n})$ given the current local model parameters $w^{t}_{n}$ in round $t$. The global optimization objective is then defined as finding the optimal set of local model parameters:
$\mathcal{L}_{n}(w^{t}_{n})=A\bigl(s_{n}^{t},a_{n}^{t}\bigr)\,\nabla_{\phi}V_{\phi}\bigl(s_{n}^{t}\bigr)+\nabla_{\theta}\log\pi_{\theta}\bigl(a_{n}^{t}\mid s_{n}^{t}\bigr)\,A\bigl(s_{n}^{t},a_{n}^{t}\bigr)$ (7)
$\bigl\{(w^{t}_{1})^{*},(w^{t}_{2})^{*}\dots,(w^{t}_{|N|})^{*}\bigr\}\;=\;\arg\!\min_{\{w^{t}_{n}\}_{n=1}^{|N|}}\;\sum_{n=1}^{|N|}\mathcal{L}_{n}(w^{t}_{n}),$ (8)
Thus, each agent strives to maximize its personalized RL objective $\mathcal{L}_{n}$, while still benefiting from collaborative updates provided by the federated framework. Formally, this MARL-adapted formulation parallels the supervised definition in the original FedFomo framework, but incorporates RL-specific objectives $\mathcal{L}_{n}$.
@@
Unlike FedAvg, which assigns equal weight to all models (or weights them by local dataset size), for each agent $n$ FedFomoLight optimizes personalized importance $\boldsymbol{\rho_{n}}=[\rho_{n1},\dots,\rho_{nN}]$, where $\sum_{i=1}^{N}\rho_{ni}=1$ to improve the local RL objective $\mathcal{L}_{n}$. According to Equ. (9), the model deltas $w_{i}^{t}-w_{n}^{t-1}$ determine how much each client should adjust its local model parameters to optimize for $\mathcal{L}_{n}$. The steps we can take toward this objective are restricted by the fixed model parameters $w_{i}^{t}$ available at time $t$. For personalized or non-IID target distributions, we can iteratively solve for the optimal combination of client models using Equ. (10) after the deviation from $t$ to $t+1$ (xiong2024personalized).
@@
$w_{n}^{t+1}\;\leftarrow\;w_{n}^{t-1}\;-\;\alpha\,\mathbf{1}^{T}\nabla_{\boldsymbol{\rho_{n}}}\;\mathcal{L}_{n}\bigl(w_{n}^{t-1}\bigr),$ (10)
@@
$\sum_{i=1}^{N}\rho_{ni}\;\bigl(w_{i}^{t}-w_{n}^{t-1}\bigr)\;=\;-\,\alpha\;\mathbf{1}^{T}\nabla_{\boldsymbol{\rho_{n}}}\;\mathcal{L}_{n}\bigl(w_{n}^{t-1}\bigr).$ (11)
To approximate $\frac{\partial}{\partial\rho_{ni}}\,\mathcal{L}_{n}(w_{n}^{t-1})$, we employ a first-order Taylor expansion. Define:
@@
Evaluating $\mathcal{L}_{n}(\varphi_{ni}(\rho_{ni}))$ near $\rho_{ni}=0$ and then setting $\rho_{ni}=1$ provides:
$\frac{\partial}{\partial\rho_{ni}}\,\mathcal{L}_{n}\bigl(w_{n}^{t-1}\bigr)\;\approx\;\mathcal{L}_{n}\bigl(w_{i}^{t}\bigr)\;-\;\mathcal{L}_{n}\bigl(w_{n}^{t-1}\bigr).$ (13)
@@
$\rho_{ni}\;=\;-\,\alpha\,\frac{\mathcal{L}_{n}\bigl(w_{i}^{t}\bigr)\;-\;\mathcal{L}_{n}\bigl(w_{n}^{t-1}\bigr)}{\bigl\|w_{i}^{t}-w_{n}^{t-1}\bigr\|}.$ (14)
@@
1:Initialize: For each agent $n\in\{1,\dots,N\}$, initialize local parameters $w_{i}^{0}$
2:Set: Number of global rounds $T$, local training steps per round $\cal T$.
3:for each global round $t=1,2,\dots,T$ do
@@
6:   Run A2C locally for $\cal T$ timesteps using $w_{n}^{t-1}$.
7:   Obtain updated local parameters $w_{n}^{t}$.
8:   Upload $w_{n}^{t}$ to the central FL server.
@@
13:   Let $\boldsymbol{\rho_{n}}=[\rho_{n1},\dots,\rho_{nN}]$ be the importance vector for agent $n$.
14:   for each agent $n$ do
15:    Compute raw importance $\rho_{ni}$ via Equ. (14)
@@
17:   Normalize to ensure $\rho_{ni}\geq 0$ and $\sum_{i}\rho_{ni}=1$
@@
21:   Update model parameters based on Equ. (16).
22:   Return updated $w_{n}^{t}$ to agent $n$.
@@ 4.3 FedClusterLight
1:Initialize: For each agent $n\in\{1,\dots,N\}$, initialize local Actor and Critic parameters $w_{n}^{0}$.
2:for each global round $t=1,2,\dots,T$ do
@@
5:   Run A2C locally for $\cal T$ timesteps using $w_{n}^{t-1}$.
6:   Obtain updated local parameters $w_{n}^{t}$.
7:   Upload $w_{n}^{t}$ to the central FL server.
@@
10:  The server groups agents into $K$ clusters $\{C_{(1)},\dots,C_{(K)}\}$ based on similarity of parameters or other relevant attributes.
@@
12:  for each cluster $C_{(k)}$ in $\{C_{(1)},\dots,C_{(K)}\}$ do
@@
17:   Determine the cluster $k$ such that $n\in C_{(k)}$.
18:   Broadcast the weights to the agents according to Equ. (19).
@@ 5.1 Experiment Set-up
We utilize Ray’s RLlib to train the proposed HFRL methods and evaluate the final policies via simulation. The road networks are shown in Fig. 3: (1) Grid 3×3 (see Fig. 3) (2) Grid 5×5 (see Fig. 3); (3) Real-world (see Fig. 3). For the 3×3 network, the primary east-west and north-south routes are designed as two-way, four-lane roads, while all other roads are two-way, two-lane roads; for the 5×5 network, the most central north-south and east-west routes are designed as two-way, six-lane roads, adjacent roads are two-way, four-lane roads, and the outermost roads are two-way, two-lane roads. The real-world network is the road networks near Columbia University generated from the OpenStreetMap [^1] https://www.openstreetmap.org, © OpenStreetMap.. For the Grid 3×3 and Grid 5×5 configurations, we designated the rightmost lane at each intersection as a mixed lane for straight-through and right-turn movements, allowing some vehicles to proceed through the intersection or turn right. The straight-through and right-turn ratios for these lanes are set at 90% and 10%, respectively. Regarding the traffic demand, the light demand is $200$ vehicles per lane per hour, while the heavy demand is $400$ vehicles per lane per hour. To build a more realistic simulation on the real-world network, we utilize two data sources to calibrate the simulation: TomTom data from the TomTom MOVE web portal  [^2] https://move.tomtom.com, © TomTom API. and manually recorded traffic counts at intersections. We then used data from the TomTom dataset to define input flows at each edge intersection, as illustrated in Fig. 4. In addition, we collected real-world traffic data from several intersections over a one-hour period, and these data will also be incorporated into our traffic flow generation file.
We employ the A2C algorithm as the underlying reinforcement learning framework for our proposed HFRL method, with the detailed hyperparameter settings summarized in Tab. 5. A2C is selected because it offers a good balance between training stability and computational efficiency. As a synchronous version of the actor–critic framework, it updates policy and value networks simultaneously, allowing the agents to learn both optimal actions and accurate value estimates. This parallel structure reduces variance in policy gradients compared with purely policy-based methods and enables training stabilization, which is particularly beneficial in multi-agent traffic signal control environments.
Table 4: Simulation Configuration for HFRL Experiments
Parameter Value / Description
Simulation software SUMO v1.19.0
Episode duration 1000 time steps
Total training episodes 100 (Synthetic networks), 50 (Real-world network)
Road networks Grid $\times3$, Grid $\times5$, Real-world
Lane configurations See Fig. 3, lane length = 100 m
Mixed lane ratio (Grid) 90% straight, 10% right-turn
Turning ratios (Real-world) $\approx$ 80% straight, 15% right-turn, 5% left-turn
Traffic demand (light) 200 veh/lane/hour
Traffic demand (heavy) 400 veh/lane/hour
Phase structure 4-phase plan: two opposing green phases with yellow transitions
Green phase duration 42 s (per direction, subject to RL control)
Yellow phase duration 3 s
Minimum green time 4 s
Maximum green time 120 s
Cluster number $K$ 4
Calibration data TomTom MOVE + manual traffic counts
Table 5: A2C Algorithm Hyperparameters
Hyperparameter Value
Framework RLlib (Ray)
Policy network architecture 2 fully connected layers, 256 units each, ReLU activation
Optimizer Adam
Learning rate $1\times 10^{-4}$
Discount factor ($\gamma$) 0.99
Entropy coefficient 0.01
Value function coefficient 0.5
Rollout length 20 steps
Batch size 4,000
Gradient clipping 40.0
Number of workers 8
Random seed 42
@@ 5.1.1 Baseline
    Centralized RL. In the centralized RL approach, a single policy network serves as a centralized controller that manages all intersections simultaneously. Each intersection periodically sends its local traffic observations to the central controller, which then processes the combined information from the entire network to determine coordinated signal actions for all intersections. Because the controller learns from global observations, it can achieve faster stabilization and more effective coordination across the network. However, this approach requires continuous communication between all intersections and the central controller, leading to significant communication overhead and high computational demand in large-scale systems.
@@ 5.1.2 Performance Metrics
    Travel Time. The travel time of vehicle $i$ is defined as
    $T_{i}=t_{i}^{\text{arrival}}-t_{i}^{\text{departure}},$
    where $t_{i}^{\text{departure}}$ and $t_{i}^{\text{arrival}}$ denote the departure and arrival times of vehicle $i$, respectively. The average travel time across all vehicles is then computed as
    $\overline{T}=\frac{1}{N_{v}}\sum_{i=1}^{N_{v}}T_{i},$
    where $N_{v}$ is the total number of vehicles. Note that in SUMO, the departure time corresponds to the time when a vehicle actually enters the network. Therefore, the computed travel time inherently includes departure delay caused by queueing, spillback, or upstream congestion, which is particularly relevant under high-demand conditions.
@@
    Waiting Time. The waiting time of vehicle $i$ is defined as the total time steps during which its speed $v_{i}(t)$ is below a small threshold, i.e.,
    $W_{i}=\sum_{t}\mathbb{I}\left[v_{i}(t)\leq 0.1\,\text{m/s}\right],$
    where $\mathbb{I}[\cdot]$ is the indicator function. The average waiting time across all vehicles is given by
    $\overline{W}=\frac{1}{N_{v}}\sum_{i=1}^{N_{v}}W_{i}.$
    Lower values of $\overline{T}$ and $\overline{W}$ indicate more efficient and coordinated traffic signal control.
@@
    Communication Cost. This is an estimated communication cost (hudson2022smart) including: (1) transmitting critic and actor parameters from edge devices to the central server; (2) distributing these parameters from the central server back to the edge devices; (3) sending actions from edge devices to traffic lights; (4) relaying observations from traffic lights to edge devices; and (5) transmitting vehicle data from vehicles to the traffic lights.
@@ 5.2.2 Performance Comparison
For the Grid 3×3 network, we present comparisons of travel time and waiting time among the proposed FedClusterLight, FedFomoLight, and baseline methods in Tab. 6. The results show that for the Grid 3×3 network under relatively low traffic demand, the centralized method outperforms both the decentralized and FedAvg methods. FedClusterLight achieves results close to those of FedAvg, while the proposed FedFomoLight method ranks second only to the centralized approach. Overall, the centralized method achieves the best performance, surpassing all other baselines. Regarding waiting time, the general trend is consistent with that of travel time.
For the Grid 5×5 network, however, the trend differs. Under light traffic conditions, as shown in Tab. 6, FedFomoLight achieves the best performance, while FedClusterLight also outperforms the other methods. In heavy traffic conditions, the centralized method performs better than the proposed methods, benefiting from complete global observations and relatively low computational and communication overhead, which allows it to coordinate all intersections more effectively.
Table 6: Travel Time and Waiting Time for Different Algorithms on Various Networks
Algorithm Volume $\times3$ Network $\times5$ Network Real-world Network
Travel $\downarrow$ Waiting $\downarrow$ Travel $\downarrow$ Waiting $\downarrow$ Travel $\downarrow$ Waiting $\downarrow$
FedFomoLight Light 53.53 10.87 $\boldsymbol{82.23}$ $\boldsymbol{17.04}$ $\boldsymbol{144.73}$ $\boldsymbol{34.20}$
Heavy 71.58 21.15 97.81 26.84 / /
FedClusterLight Light 57.90 12.85 83.44 17.81 148.70 36.62
Heavy 69.65 19.6 100.22 30.46 / /
FedAvg Light 58.57 13.42 85.77 18.36 157.41 37.16
Heavy 77.85 23.35 105.61 31.04 / /
Decentralized Light 57.18 12.46 87.23 23.40 643.75 501.38
Heavy 78.31 28.25 112.11 66.45 / /
Centralized Light $\boldsymbol{50.11}$ $\boldsymbol{8.03}$ 83.20 17.62 356.01 247.95
Heavy $\boldsymbol{69.32}$ $\boldsymbol{19.53}$ $\boldsymbol{96.08}$ $\boldsymbol{24.85}$ / /
MPlight Light 62.39 80.41 119.83 30.60 710.27 210.09
Heavy 73.35 47.83 138.54 38.85 / /
Fixed time Light 86.75 17.37 150.19 96.11 767.12 311.05
Heavy 103.55 78.45 209.25 159.34 / /
Actuated Light 75.24 26.09 142.32 55.02 548.35 138.29
Heavy 99.45 34.57 176.83 100.98 / /
We compare communication costs across different traffic networks under various traffic demand scenarios in Tab. 7 and Tab. 8. The centralized method typically requires uploading all local observations to a central server for computation, which leads to significantly higher communication costs. In contrast, FRL and the decentralized method generally maintain similarly low communication costs. FRL methods, due to the allowance for model information exchange, ensure more stable communication costs during training.
@@
We also report the average training time per epoch in Tab. 9. As shown in the table, the training time for distributed methods does not differ significantly from that of the centralized approach. In fact, the proposed methods achieve slightly lower training times due to their improved performance, which accelerates the stabilization of simulation trials. To further examine scalability, we conducted additional experiments on larger networks with 100 and 150 intersections on FedFomoLight. The average training time per epoch increases from approximately 550 s to 689.15 s as the number of intersections grows. As shown in Fig. 7, the training time increases approximately linearly with the number of intersections. This demonstrates that the proposed HFRL framework can scale to large networks while maintaining reasonable computational efficiency. Although communication and synchronization overhead inevitably grow with network size, the hierarchical grouping in FedFomoLight and FedClusterLight effectively mitigates these issues by limiting parameter aggregation within smaller clusters, rather than across all agents globally.
Figure 7: Scalability of Training Time with Increasing Number of Intersections
Table 9: Average Training Time on each Epoch.
Training Time Each Epoch (Seconds)
Methods Grid 3x3 Grid 5x5 Cosmos
FedFomoLight 71.8797 97.6731 156.6420
FedClusterLight 72.3310 99.0668 156.6511
FedAvg 71.3830 101.1719 159.7703
Decentralized 74.0686 99.8479 159.9270
Centralized 59.6006 87.8595 145.9587